# Parallel Split Learning with Global Sampling

### v4

@@ IV-B Deviation Analysis
Although the $\{Y_{m}\}^{m\in\mathcal{M}}$ are dependent as they are required to sum up to $B$, a union bound yields
$\Pr\!\Bigl{(}\Delta(\mathcal{B},\boldsymbol{\beta}_{0})\geq M\varepsilon\Bigr{)}$ $\leq\sum_{m\in\mathcal{M}}2\exp\!\Bigl{(}-\tfrac{2\,\varepsilon^{2}\,B}{1-\tfrac{B-1}{D_{0}}}\Bigr{)}$ (9)
$=2M\exp\!\Bigl{(}-\tfrac{2\,\varepsilon^{2}\,B}{1-\tfrac{B-1}{D_{0}}}\Bigr{)}.$ (10)
By construction, GPSL’s depletion‐aware sampling is exactly equivalent to drawing $B$ points uniformly at random without replacement from the pooled dataset $\mathcal{D}_{0}$. Consequently, the class counts $Y_{m}$ in GPSL follow the same hypergeometric law as in CL. Applying Serfling’s inequality with finite‐population correction and then a union bound over $\mathcal{M}$ gives
$\Pr\!\Bigl{(}\Delta(\mathcal{B},\boldsymbol{\beta}_{0})\geq M\varepsilon\Bigr{)}$ $=\sum_{m\in\mathcal{M}}\Pr\!\Bigl{(}\bigl{|}\tfrac{Y_{m}}{B}-\beta_{0,m}\bigr{|}\geq\varepsilon\Bigr{)}$ (11)
$\leq 2M\exp\!\Bigl{(}-\frac{2\,\varepsilon^{2}\,B}{1-\tfrac{B-1}{D_{0}}}\Bigr{)}.$ (12)
Thus GPSL attains the exact same exponential‐tail deviation guarantee (up to the standard finite‐population factor) as centralized uniform sampling.
IV-C Comparison with Fixed Local Batches
Under a scheme with a fixed local batch size as in [8], each client $k$ contributes $\tilde{B}_{k}=\lceil B\cdot\tfrac{D_{k}}{D_{0}}\rceil$ samples without replacement in each step. Defining $\tilde{Y}^{\prime}_{m}=\sum_{k\in\mathcal{K}}\tilde{Y}_{k,m}$ with
and $\tilde{p}_{m}=\sum_{k\in\mathcal{K}}\tfrac{\tilde{B}_{k}}{B}\,\beta_{k,m}$ we decompose
The sampling noise term still decays exponentially, but the rounding bias satisfies
$\bigl{|}\tilde{p}_{m}-\beta_{0,m}\bigr{|}$ $=\Bigl{|}\sum_{k\in\mathcal{K}}\Bigl{(}\tfrac{\tilde{B}_{k}}{B}-\tfrac{D_{k}}{D_{0}}\Bigr{)}\beta_{k,m}\Bigr{|}$ (14)
$\leq\;\sum_{k\in\mathcal{K}}\Bigl{|}\tfrac{\tilde{B}_{k}}{B}-\tfrac{D_{k}}{D_{0}}\Bigr{|}=O\!\bigl{(}\tfrac{K}{B}\bigr{)}.$ (15)
When $K/B$ is large, the rounding bias can dominate and renders any exponential‐tail guarantee ineffective. By contrast, GPSL’s server‐driven batch allocation incurs zero per‐class rounding bias while retaining the same exponential decay. Fixed local batches may still perform adequately when $K\ll B$ and rounding effects are negligible. However, such settings are rare in edge learning environments, where large $K$ and small batch sizes dominate due to memory and communication constraints.
Moreover, the rounding exacerbates the issue of client data depletion, where client datasets are depleted disproportionately and faster. This increases the total number of training steps $T$ and prolongs training. By contrast, GPSL prevents unnecessary inflation of $T$ by keeping the global batch size $B$ constant, regardless of the number of clients.
V Simulation Results
[…]
V-A Sampling Methods
@@
In Table III we report the results of the sampling methods in the IID setting, where all clients have the same distribution. All sampling methods perform similarly, achieving near central learning performance.
V-B Non-IID Data
In the non-IID scenario, we employ a modified version of the extended Dirichlet allocation method proposed in [7], which ensures that each client receives data from exactly $C$ classes. First, $C$ distinct classes are assigned to each client using a multinomial-based allocation that promotes balanced class coverage across the client population. Then, for each class, a Dirichlet distribution with concentration parameter $\alpha$ is sampled to determine how its samples are distributed among the clients assigned to it. Lower values of $\alpha$ lead to more imbalanced allocations, where a class is concentrated heavily on one or a few clients. Higher values of $\alpha$ result in more uniform class sample distributions across clients. By tuning $C$ and $\alpha$, we can control the degree and structure of statistical heterogeneity across clients. We consider a mild non-IID scenario with $C=5$ and $\alpha=3.0$ and a severe non-IID scenario with $C=2$ and $\alpha=3.0$. In Figure 2 we show an example of the data distribution for $K=16$ clients in the severe non-IID scenario.
Figure 2: Example of a severe non-IID setting for $K=16$ clients, showing the distribution of data across classes.
TABLE I: IID test accuracy on CIFAR10
@@ V-B Non-IID Data
TABLE II: Severe Non-IID test accuracy on CIFAR10
@@
TABLE III: Mild Non-IID test accuracy on CIFAR10
@@
Top-1 test accuracy on CIFAR-10 under IID and non-IID settings for different numbers of clients $K$ and global batch size $B=128$. Each result is averaged over 5 random seeds and reported with standard deviation. CL achieved 84.06±0.5384.06$\pm$ 0.5384.06 ± 0.53. Bold values indicate the best result per column.
Figure 3: Test accuracy curves for each sampling method over 5 runs with standard deviation shaded. Severe non-IID setting with $K=64$ clients and a global batch size of $B=128$.
Table III shows the performance of each sampling method under severe non-IID conditions. Table III shows the performance in a mild non-IID scenario. In both cases, GPSL achieves a test accuracy on par with centralized learning, while FPLS and FLS degrade significantly. The results are consistent across both mild and severe non-IID scenarios for varying client counts $K$, indicating scalability and robustness to the degree of data heterogeneity. Figure 3 shows the test accuracy curves for each sampling method under severe non-IID conditions. GPSL maintains a stable convergence curve, while FPLS and FLS exhibit significant fluctuations.
TABLE IV: Effect of Global Batch Size on Severe Non-IID Test Accuracy
@@
Top-1 test accuracy on CIFAR-10 under severe non-IID settings for different global batch sizes $B$. Each result is averaged over 5 random seeds and reported with standard deviation. $K=128$ clients. Bold values indicate the best result per column.
V-C Global Batch Size
Table IV shows performance results for the different sampling methods with varying batch sizes $B$. GPSL consistently outperforms FPLS and FLS across all batch sizes, demonstrating its robustness to batch size variations. In particular, GPSL enables flexibility in choosing the global batch size whereas FPLS and FLS are sensitive to the choice of batch size and may result in much larger global batch sizes than intended due to rounding and the client data depletion problem (Section IV-C).
V-D Batch Deviation
Figure 4: Batch deviation curves for each sampling method over 5 runs with exponential moving average smoothing. Severe non-IID setting with $K=64$ clients and a global batch size of $B=128$.
Figure 4 shows the empirical evaluation of the batch deviation as defined in Section IV-B for each sampling method under severe non-IID conditions. GPSL maintains a low and stable batch deviation almost matching CL, while FPLS and FLS exhibit significant fluctuations, particularly in the standard deviation. This aligns with our theoretical analysis and performance results, indicating that the batch deviation is the main reason for the performance degradation of FPLS and FLS.
@@ V-D Batch Deviation
V-E Runtime
Figure 5: Average total training time for each sampling method over 5 runs, measured in minutes. Severe non-IID setting with $K=64$ clients.
@@ V-E Runtime
V-F Limitations
While GPSL improves model performance and training efficiency, it inherits the PSL framework’s reliance on gradient averaging, which increases training latency and communication overhead. This limits its applicability to real-time and unstable IoT environments. Future work should explore integrating GPSL with client clustering [27] or communication-efficient techniques, such as adaptive compression [10], to improve scalability and reduce overhead.
@@ VI Conclusion
We introduced parallel split learning with global sampling, a method designed to address the large effective batch size and non-independent and identically distributed data problems in distributed deep learning. By dynamically adjusting local batch sizes while maintaining a fixed global batch size, the method decouples the global batch size from the number of participating clients. Using standard concentration bounds, we showed that global sampling improves the alignment between global batches and the overall data distribution. Our simulation results confirm that the proposed method enhances convergence stability, boosts classification accuracy, and reduces training latency in non-independent and identically distributed scenarios. These findings suggest that server-driven global sampling offers a simple yet effective improvement for scalable distributed learning at the network edge.

### v5

@@ V-A Experimental Setup
TABLE I: Experimental setup.
Aspect Configuration
Dataset CIFAR-10: 50k train / 10k test, 10 classes
Splits IID; Non-IID (mild: $C{=}5,\,\alpha{=}3.0$), (severe: $C{=}2,\,\alpha{=}3.0$)
Clients $K\in\{16,32,64,128\}$ (default K=64K{=}64)
Global batch $B\in\{64,128,256\}$ (default B=128B{=}128)
Cut layer Residual layer 3 (server)
Optimizer SGD with momentum $0.9$, weight decay $5\times 10^{-4}$, learning rate $10^{-2}$
Loss Cross entropy with label smoothing ($\epsilon=0.1$)
Epochs E=100E{=}100
Augmentation Random crop, random horizontal flip
V-B Sampling Methods
@@
In Table IV we report the results of the sampling methods in the IID setting, where all clients have the same distribution. All sampling methods perform similarly, achieving performance on par with the centralized learning (CL) baseline.
V-C Non-IID Data
In the non-IID scenario, we employ a modified version of the extended Dirichlet allocation method proposed in [7], which ensures that each client receives data from exactly $C$ classes. First, $C$ distinct classes are assigned to each client using a multinomial-based allocation that promotes balanced class coverage across the client population. Then, for each class, a Dirichlet distribution with concentration parameter $\alpha$ is sampled to determine how its samples are distributed among the clients assigned to it. Lower values of $\alpha$ lead to more imbalanced allocations, where a class is concentrated heavily on one or a few clients. Higher values of $\alpha$ result in more uniform class sample distributions across clients. By tuning $C$ and $\alpha$, we can control the degree and structure of statistical heterogeneity across clients. Thus, $C$ governs the per-client label diversity (a hard cap on how many distinct classes a client can observe), while $\alpha$ controls the smoothness of intra-class sharing across the eligible clients. The pair $(C,\alpha)$ therefore parameterizes the severity and structure of the data heterogeneity. We consider a mild non-IID scenario with $C=5$ and $\alpha=3.0$ and a severe non-IID scenario with $C=2$ and $\alpha=3.0$. In Figure 2 we show an example of the class distribution for $K=16$ clients in the severe non-IID scenario.
Figure 2: Example of a severe non-IID setting for $K=16$ clients, showing the distribution of classes across clients.
TABLE II: IID Test Accuracy
@@ V-C Non-IID Data
TABLE III: Severe non-IID Test Accuracy $(C=2,\alpha=3.0)$
@@
TABLE IV: Mild non-IID Test Accuracy $(C=5,\alpha=3.0)$
@@
ResNet-18 on CIFAR-10: Top-1 test accuracy under IID and non-IID settings for different numbers of clients $K$. CL achieved $84.06\,\pm\,0.53$. Bold indicates the best per column.
Figure 3: ResNet-18 on CIFAR-10: Test accuracy curves for each sampling method in a severe non-IID setting with standard deviation shaded.
Table IV shows the performance of each sampling method under severe non-IID conditions. Table IV shows the performance in a mild non-IID scenario. In both cases, GPSL achieves a test accuracy on par with centralized learning, while FPLS and FLS degrade significantly. The results are consistent across both mild and severe non-IID scenarios for varying client counts $K$, indicating scalability and robustness to the degree of data heterogeneity. Figure 3 shows the test accuracy curves for each sampling method under severe non-IID conditions. GPSL maintains a stable convergence curve, while FPLS and FLS exhibit significant fluctuations.
TABLE V: Effect of global batch size on test accuracy
@@
ResNet-18 on CIFAR-10: Top-1 test accuracy under severe non-IID setting for different global batch sizes $B$ with $K=128$ clients. Bold indicates the best per column.
V-D Global Batch Size
Table V shows performance results for the different sampling methods with varying global batch sizes $B$. GPSL consistently outperforms FPLS and FLS across all batch sizes, demonstrating its robustness to batch size variations. In particular, GPSL enables flexibility in choosing the global batch size whereas FPLS and FLS are sensitive to the choice of batch size and may result in much larger global batch sizes than intended due to rounding and the client data depletion problem (Section IV-C).
V-E Batch Deviation
Figure 4: ResNet-18 on CIFAR-10: Batch deviation curves for each sampling method in a severe non-IID setting with exponential moving average smoothing and shaded standard deviation.
Figure 4 shows the empirical evaluation of the batch deviation per-step as defined in Section IV-B for each sampling method under severe non-IID conditions. GPSL maintains a low and stable batch deviation almost matching CL, while FPLS and FLS exhibit significant fluctuations, particularly in the standard deviation. This aligns with our theoretical analysis and performance results, indicating that the batch deviation is the main reason for the performance degradation of FPLS and FLS.
@@ V-E Batch Deviation
V-F Runtime
Figure 5: ResNet-18 on CIFAR-10: Average total training time for each sampling method in a severe non-IID setting, measured in minutes.
@@ V-F Runtime
V-G Limitations
In settings with $K>B$, some clients necessarily receive zero allocations in a step, reducing instantaneous utilization and potentially inflating wall-clock time. Lightweight grouping or clustering of clients can mitigate this idle capacity [18]. Our experiments assume a static, synchronous, straggler-free population, whereas real deployments must handle churn, delayed or dropped contributions, and network jitter. Because GPSL only changes how samples are allocated, established methods, such as asynchronous scheduling or deadline/slack participation remain directly compatible to address these systems issues [28]. However, designing novel approaches for dealing with such instable situations specific to PSL remains an open research direction. Another direction for future work is to integrate churn-aware re-sampling and adaptive policies that jointly tune the global sampling and participation pattern in highly dynamic environments such as edge and IoT deployments.
@@ VI Conclusion
We introduced parallel split learning with global sampling (GPSL), a drop‑in sampling mechanism that fixes the global batch size and allocates per‑client batch size schedules using pooled‑level proportions while clients sample locally. By eliminating per‑class rounding and decoupling the effective batch size from the number of participating clients, GPSL recovers the statistics of centralized uniform sampling without replacement and admits finite‑population deviation guarantees. Empirically, under non‑IID splits GPSL consistently attains centralized‑like accuracy and stabilizes optimization. Furthermore, GPSL shortens training time by avoiding inflation of steps induced by client data depletion without increasing server memory or coordination overhead. Relative to fixed local batching, GPSL changes only how batches are formed while preserving the PSL training loop, thereby removing the additive bias that skews global label histograms and inflates the effective batch size. The main limitations are scope and setting: our evaluation assumes a static, synchronous population. When $K>B$, per‑step idleness may limit utilization. These constraints are orthogonal to GPSL and can be addressed with existing PSL techniques. Looking ahead, we see value in churn‑aware re‑sampling and adaptive policies that jointly tune global batch size and participation. Collectively, these properties make GPSL practical for edge and IoT deployments, where resources are limited and client populations are large and heterogeneous
Acknowledgment
This work was funded by the German Research Foundation (DFG) under the Cluster of Excellence CARE: Climate- Neutral And Resource-Efficient Construction (EXC 3115), project number 533767731 and by the Federal Ministry of Re- search, Technology, and Space (BMFTR, Germany) as part of NeuroSys: Efficient AI-methods for neuromorphic computing in practice (Projekt D) - under Grant 03ZU2106DA.
The authors used language-assistance tools for copy-editing; responsibility for the content rests with the authors.
-A Additional Results on CIFAR-100 and ResNet-34
We provide additional experiments on the CIFAR-100 dataset [24] and a deeper ResNet-34 backbone [25]. All experimental settings follow Section V except for a longer training horizon of E=200E{=}200 epochs. Results are averaged over 3 runs and reported with standard deviation.
Table VI summarizes results in the severe non-IID case ($C{=}2,\alpha{=}3.0$) with K=64K{=}64 and B=128B{=}128. GPSL again matches CL on both datasets, while FLS and FPLS trail, with a larger gap on CIFAR-100. In Figure 6, GPSL’s test accuracy on CIFAR-100 rises smoothly to a higher plateau with tight variability, whereas FLS and FPLS settle lower and fluctuate more. Figure 7 shows GPSL’s batch deviation closely tracking CL, while FLS and FPLS remain higher and more volatile. Notably, relative to the CIFAR-10/ResNet-18 experiment, FLS exhibits a lower batch deviation yet still shows a pronounced accuracy gap. This reflects increased task difficulty and model depth: small but persistent distributional mismatch spread across many classes over a longer training horizon can impede representation learning, so lower absolute deviation values need not translate into comparable accuracy. Collectively, these observations indicate GPSL’s suitability as a robust drop-in choice across architectures and datasets.
TABLE VI: Severe Non-IID Test Accuracy
Method CIFAR-10 CIFAR-100
FLS 25.42 $\pm$ 7.42 41.61 $\pm$ 2.88
FPLS 13.62 $\pm$ 2.75 3.13 $\pm$ 1.24
GPSL 89.69 $\pm$ 0.24 60.04 $\pm$ 1.18
ResNet-34 on CIFAR-10 and CIFAR-100: Top-1 test accuracy under a severe non-IID setting. CL achieved $89.62\pm 1.16$ on CIFAR-10 and $59.86\pm 1.45$ on CIFAR-100. Bold indicates the best per column.
Figure 6: ResNet-34 on CIFAR-100: Test accuracy curves for each sampling method in a severe non-IID setting with standard deviation shaded.
Figure 7: ResNet-34 on CIFAR-100: Batch deviation curves for each sampling method in a severe non-IID setting with exponential moving average smoothing and shaded standard deviation.